# Digging into contrastive learning for robust depth estimation with diffusion models

### v3

@@ Abstract.
Recently, diffusion-based depth estimation methods have drawn widespread attention due to their elegant denoising patterns and promising performance. However, they are typically unreliable under adverse conditions prevalent in real-world scenarios, such as rainy, snowy, etc. In this paper, we propose a novel robust depth estimation method called D4RD, featuring a custom contrastive learning mode tailored for diffusion models to mitigate performance degradation in complex environments. Concretely, we integrate the strength of knowledge distillation into contrastive learning, building the ‘trinity’ contrastive scheme. This scheme utilizes the sampled noise of the forward diffusion process as a natural reference, guiding the predicted noise in diverse scenes toward a more stable and precise optimum. Moreover, we extend noise-level trinity to encompass more generic feature and image levels, establishing a multi-level contrast to distribute the burden of robust perception across the overall network. Before addressing complex scenarios, we enhance the stability of the baseline diffusion model with three straightforward yet effective improvements, which facilitate convergence and remove depth outliers. Extensive experiments demonstrate that D4RD surpasses existing state-of-the-art solutions on synthetic corruption datasets and real-world weather conditions. The code for D4RD will be made available for further exploration and adoption.
@@
copyright: acmcopyrightjournalyear: 2022doi: XXXXXXX.XXXXXXXsubmissionid: 2846conference: Conference; Some Days; Anywhere on earthprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reconstruction
@@ 2.2. Robust Depth Estimation
Contrastive-based methods typically improve the model’s robustness by ensuring consistency between the depth predictions of a clear image and its augmented versions. Robust-Depth(Saunders et al., 2023) introduces the semi-augmented warp and bi-directional contrast, which ensures the consistency assumption and improve the accuracy. WeatherDepth considers the domain gap between clear scenes and complex scenes and solves this problem by applying a gradual adaptation scheme based on curriculum contrastive learning. Additionally, EC-Depth(Song et al., 2023) combines contrast and distillation with double-stage training. The first stage uses contrast, creating a KL divergence-like contrast to align the depth of harsh scenes, and the second stage uses distillation, introducing EMA for teacher-student joint learning.
@@ 3. Method
Figure 2. The training framework of D4RD is depicted at the upper part of the image. Below that, multi-level trinity learning (i.e. images, depth features, and noise prediction) are presented through (a), (b), and (c), respectively.
@@ 3.1. Preliminaries
(2) $L_{ph}=\omega_{1}\frac{1-\operatorname{SSIM}\left(\textbf{I},\textbf{I}_{a\rightarrow t}\right)}{2}+\omega_{2}\left\|\textbf{I}-\textbf{I}_{a\rightarrow t}\right\|.$
@@
(3) $\textbf{I}_{a\rightarrow t}^{\prime}=\textbf{I}_{a}\left\langle\operatorname{proj}\left(\textbf{d}_{aug},T_{t\rightarrow a},K\right)\right\rangle,$
@@
Obviously, if $\epsilon_{\theta}$ is completely consistent with the $\epsilon$, we can reverse the depth result precisely, so we have the classical training loss:
(6) $L_{ddim}=\mathbb{E}_{\textbf{d}_{0},\epsilon\sim\mathcal{N}(0,1),\tau\sim\mathcal{U}(T)}\left\|\epsilon-{\epsilon_{\theta}}\left({{\textbf{d}_{\tau-1}}\left|{{\textbf{d}_{\tau}},\textbf{{c}}}\right.}\right)\right\|^{2}_{2},$
where $\mathcal{U}(\cdot)$ is the uniformly sampling. For self-supervised learning, D4RD adopts the pseudo-GT diffusion as used in MonoDiffusion (Shao et al., 2023), which introduces a pre-trained model $\mathcal{F}_{T}$ to infer the pseudo-depth $\textbf{d}_{T}$ and it will replace $D$ in Eq. 4.
@@ 3.2. Stable Depth with Diffusion Models
where $\mathcal{V}(\cdot)$ is a pre-trained variational-auto-encoder (VAE) and $\mathcal{V}^{-1}(\cdot)$ is the reverse decoder. This manner undoubtedly introduces extra calculations. In this paper, we simplify the above solution through a non-linear activation function (and its inverse):
@@
where $\oplus$ represents the channel-dim concatenation. We visualize the feature map in Fig. 4 to analyze the reasons and reveal that $\mathcal{F}(\textbf{I})$ is a coarse ‘depth map’ with some additional incorrect edges. In other words, $\mathcal{F}_{f}$ focuses solely on depth-relevant information, and the combination in Eq. 11 provides higher-level contextual and semantic information, which helps the diffusion architecture in understanding the scenes, predicting noises, and ultimately generating better estimations.
@@ 3.3. Trinity Contrastive learning
where $\underline{\text{underline}}$ denotes the predicted from a frozen model. However, Fig. 3(a) has shown that this strategy struggles to teach $\mathcal{F}_{S}$ when the augmented image $\textbf{I}_{aug}$ has a significant domain difference. Moreover, there is an evident performance upper bound for $\mathcal{F}_{S}$.
@@
In diffusion models, we discovered that in Eq. 4, the added noise $\epsilon$, which is sampled from $\mathcal{N}\left({0,I}\right)$, can serve as a perfect guiding anchor for the estimated noise $\epsilon_{\theta}$ in Eq. 6 and do not require any extra annotation. Let $\epsilon_{\theta}$ and $\epsilon_{\theta,\text{aug }}$ represent the noise prediction for the clear image I and augment image $\textbf{I}_{aug}$. Once these noises are identical and accurate, the final denoising depth will demonstrate consistency and robustness. Hence, as shown in Fig. 2(c), we reformulate a novel noise-level trinity contrast scheme as:
@@ 3.4. Multi-level Contrast
Figure 4. Visual results of each level trinity. Compared to the origin image, as shown in the highlight regions in the red dashed rectangle, the image level trinity can assist in handling water surface artifacts and ground snow. The feature level trinity builds a coarse depth map with some wrong edges but the noise trinity fixes them all. Better viewed when zooming in.
@@
Feature level. We first extend it to the depth feature level, which exploits the robust ability of feature network $\mathcal{F}$. However, at this time, we lack the guidance as perfect as the added noise. Instead, as depicted in Fig. 2 (b), we adopt a teacher model $\mathcal{F}_{T}$ to get the suboptimal guiding label and build the feature-level ‘trinity’ loss as:
(15) $\begin{array}[]{l}L_{feat}=\omega_{1}\|\mathcal{F}_{S}(\textbf{I})-\mathcal{F}_{S}(\textbf{I}_{aug})\|_{1}+\\ \omega_{2}(\|\mathcal{F}_{T}(\textbf{I})-\mathcal{F}_{S}(\textbf{I})\|_{1}+\|\mathcal{F}_{T}(\textbf{I})-\mathcal{F}_{S}(\textbf{I}_{aug})\|_{1}),\end{array}$
@@
(17) $L_{img}=\beta_{1}(\|\textbf{I}_{\text{aug }}^{\prime}-\textbf{I}\|_{1}+\|\textbf{I}^{\prime}-\textbf{I}\|_{1})+\beta_{2}\|\textbf{I}_{\text{aug }}^{\prime}-\textbf{I}^{\prime}\|_{1},$
where the values of $\beta_{1}$ and $\beta_{2}$ are same as those of $\omega_{1}$ and $\omega_{2}$.
@@
Table 1. Quantitative results on WeatherKITTI and KITTI dataset. For the error-based metrics, the lower value is better; and for the accuracy-based metrics, the higher value is better. The best and second best results are marked in bold and underline.
@@
Figure 5. Qualitative results for WeatherKITTI dataset. We compare D4RD with the current SoTA RMDE methods in the adverse rain and snow subsets. The part marked with ‘Clear’ is the correspondence sunny image (processed for more clarity). Regions with prominent differences are highlighted using dashed boxes.
@@ 3.5. Two stage training
The second training phase. Building upon the optimization objective of stage one, we add the $L_{feat}$ and $L_{img}$ to build the multi-level contrasts as mentioned in Section 3.4. The total loss of the second stage can be written as:
(19) $L_{stage2}=L_{nis}+L_{dis}+L_{ph}+\rho L_{e}+\theta(L_{feat}+L_{img}),$
where $\theta$ is empirically set to 0.5.
Figure 6. Qualitative results for real weather dataset. Besides the RMDE methods, we also compare D4RD with the standard method MonoViT(Zhao et al., 2022b) in snowy and rainy environments to show our method’s practicality. The part marked with Magnification is processed for more clarity.
@@ 4.1. Dataset
WeatherKITTI(Wang et al., 2024) is built on the KITTI dataset and includes 6 kinds of weather augmentation. We train D4RD on this dataset because, compared with other synthetic datasets(Saunders et al., 2023; Mao et al., 2024), it is more realistic in simulating real-world complex scenarios. WeatherKITTI shares the same train/validate/test splits as KITTI. Therefore, for training, we follow Zhou’s split(Zhou et al., 2017) (containing 19,905 training images and 2,212 validation images), and for testing, we adopt Eigen’s spilt(Eigen et al., 2014) with 697 test images.
KITTI-C(Kong et al., 2023) is a synthetic dataset building for the RoboDepth Competition. In total, the benchmark contains 18 types of perturbations with 5 levels of corruption magnitudes, most of which are stronger than WeatherKITTI. We use it to fairly compare D4RD with their competition champions(Song et al., 2023) to further demonstrate our robust performance.
@@
Dense(Bijelic et al., 2020) is one of the few real snowy datasets and it has approximately 1,500 snowy images in total. We introduce this benchmark to test the snowy scenes. To ensure consistency with the above dataset, we followed (Wang et al., 2024) and sampled 1 image every 3 to obtain 500 images as test data. To generate accurate depth GT in challenging snowy conditions, we use the DROR(Charron et al., 2018) algorithm to remove inaccurate depth readings from LiDARs affected by snowflakes. In addition, a small portion of distorted areas has been trimmed, resulting in a final resolution of $1880\times 924$. We will explain the reasons for not using the snowy part of CADC dataset in the supplementary materials.
@@ 4.2. Implementation Details
D4RD is implemented in PyTorch and trained on a single NVIDIA RTX 3090. We use the AdamW optimizer(Loshchilov and Hutter, 2019) and other similar hyperparameters as MonoViT. The batch size is set to 6, and the training process runs 30 epochs in each stage. Considering the resolution difference across the datasets, we resize them to $640\times 192$ for both training and testing. For the diffusion process, we follow MonoDiffusion(Shao et al., 2023) and set it with 1000 training steps and 20 inference steps.
For fair comparisons, we retrained Robost-Depth(Saunders et al., 2023) and EC-Depth(Song et al., 2023) on the WeatherKITTI dataset. Furthermore, to demonstrate D4RD’s ability to handle robust scenes, we also retrain our model with the same training data used by (Song et al., 2023), named D4RD, to compare the performance on the KITTI-C dataset.
@@ 4.3. Quantitative and Qualitative Comparison
Here, we adopt the same evaluation metrics and errors up to 80m for all comparisons as (Godard et al., 2019). Please refer to the supplementary for more details. The comparative methods include the baseline MonoDiffusion, traditional RMDE models’ baseline MonoViT(Zhao et al., 2022b), previous SoTA RMDE models WeatherDepth(Saunders et al., 2023), EC-Depth(Song et al., 2023), Robust-Depth(Saunders et al., 2023), and their retained models. Besides, (Wang et al., 2024) and (Wang et al., 2024) apply their approaches to various baselines, we want to clarify that we choose the one that uses MonoViT(Zhao et al., 2022b). In (Song et al., 2023), they take a two-stage training pipeline and we select the second stage model for the comparison.

### v4

@@ Abstract.
Recently, diffusion-based depth estimation methods have drawn widespread attention due to their elegant denoising patterns and promising performance. However, they are typically unreliable under adverse conditions prevalent in real-world scenarios, such as rainy, snowy, etc. In this paper, we propose a novel robust depth estimation method called D4RD, featuring a custom contrastive learning mode tailored for diffusion models to mitigate performance degradation in complex environments. Concretely, we integrate the strength of knowledge distillation into contrastive learning, building the ‘trinity’ contrastive scheme. This scheme utilizes the sampled noise of the forward diffusion process as a natural reference, guiding the predicted noise in diverse scenes toward a more stable and precise optimum. Moreover, we extend noise-level trinity to encompass more generic feature and image levels, establishing a multi-level contrast to distribute the burden of robust perception across the overall network. Before addressing complex scenarios, we enhance the stability of the baseline diffusion model with three straightforward yet effective improvements, which facilitate convergence and remove depth outliers. Extensive experiments demonstrate that D4RD surpasses existing state-of-the-art solutions on synthetic corruption datasets and real-world weather conditions. Source code and data are available at https://github.com/wangjiyuan9/D4RD.
@@
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681168isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Reconstruction
@@ 2.2. Robust Depth Estimation
Contrastive-based methods typically improve the model’s robustness by ensuring consistency between the depth predictions of a clear image and its augmented versions. Robust-Depth(Saunders et al., 2023) introduces the semi-augmented warp and bi-directional contrast, which ensures the consistency assumption and improves the accuracy. WeatherDepth considers the domain gap between clear scenes and complex scenes and solves this problem by applying a gradual adaptation scheme based on curriculum contrastive learning. Additionally, EC-Depth(Song et al., 2023) combines contrast and distillation with double-stage training. The first stage uses contrast, creating a KL divergence-like contrast to align the depth of harsh scenes, and the second stage uses distillation, introducing EMA for teacher-student joint learning.
@@ 3. Method
Figure 2. The training framework of D4RD is depicted in the upper part of the image, where the teacher model $\mathcal{F}_{T}$ is enclosed in the green dashed box, while the student network $\mathcal{F}_{S}$ comprises the remaining parts. There are only 3 components in the whole network: the base feature network (having two symbols, $\mathcal{F}_{fT}$ in $\mathcal{F}_{T}$, $\mathcal{F}_{fS}$ in $\mathcal{F}_{S}$), the diffusion process(grey dashed box), and the robust CNN $\mathcal{F}_{R}$. Below that, multi-level trinity learning (i.e., images, depth features, and noise prediction) is presented through (a), (b), and (c), respectively.
@@ 3.1. Preliminaries
To clarify the method, we use abundant symbols. A table of symbol meanings is provided in the supplementary materials for reference.
@@
Afterwards, we can compute the photometric reconstruction loss between I and $\textbf{I}_{a\rightarrow t}$ to constrain the depth:
(2) $L_{ph}=\varepsilon_{1}\frac{1-\operatorname{SSIM}\left(\textbf{I},\textbf{I}_{a\rightarrow t}\right)}{2}+\varepsilon_{2}\left\|\textbf{I}-\textbf{I}_{a\rightarrow t}\right\|.$
@@
(3) $\textbf{I}_{a\rightarrow t}^{\prime}=\textbf{I}_{a}\left\langle\operatorname{proj}\left(\textbf{d}_{aug},\mathcal{T}_{t\rightarrow a},K\right)\right\rangle,$
@@
Obviously, if $\epsilon_{\theta}$ is completely consistent with the $\epsilon$, we can reverse the depth result precisely, so we have the training loss:
(6) $L_{ddim}=\mathbb{E}_{\textbf{d}_{0},\epsilon\sim\mathcal{N}(0,1),\tau\sim\mathcal{U}(\mathbb{T})}\left\|\epsilon-{\epsilon_{\theta}}\left({{\textbf{d}_{\tau-1}}\left|{{\textbf{d}_{\tau}},\textbf{{c}}}\right.}\right)\right\|^{2}_{2},$
where $\mathcal{U}(\cdot)$ represents uniformly sampling. For self-supervised learning, D4RD adopts the pseudo-GT diffusion as used in MonoDiffusion (Shao et al., 2023), which introduces a pre-trained model $\mathcal{F}_{T}$ to infer the pseudo-depth $\textbf{d}_{T}$ and it will replace $D$ in Eq. 4.
@@ 3.2. Stable Depth with Diffusion Models
where $\mathcal{V}(\cdot)$ is a pre-trained variational-auto-encoder (VAE), $\mathcal{V}^{-1}(\cdot)$ is the reverse decoder, and $\textbf{d}_{0}$ no longer represents the final prediction. This manner undoubtedly introduces extra calculations. In this paper, we simplify the encoder and decoder as sigmoid-1 and sigmoid, respectively:
@@
where $\oplus$ represents the channel-dim concatenation. We visualize the feature map in Fig. 4 to analyze the reasons and reveal that $\mathcal{F}_{f}(\textbf{I})$ is a coarse ‘depth map’ with some additional incorrect edges. In other words, $\mathcal{F}_{f}$ focuses solely on depth-relevant information, and the combination in Eq. 11 provides higher-level contextual and semantic information, which helps the Fdiffusion architecture in understanding the scenes, predicting noises, and ultimately generating better estimations.
@@ 3.3. Trinity Contrastive learning
where $\underline{\text{underline}}$ denotes the prediction from a frozen model. However, Fig. 3(a) has shown that this strategy struggles to teach $\mathcal{F}_{S}$ when the augmented image $\textbf{I}_{aug}$ has a significant domain difference. Moreover, there is an evident performance upper bound for $\mathcal{F}_{S}$.
@@
In diffusion models, we discovered that in Eq. 4, the added noise $\epsilon$, which is sampled from $\mathcal{N}\left({0,I}\right)$, can serve as a perfect guiding anchor for the estimated noise $\epsilon_{\theta}$ in Eq. 6 and do not require any extra annotation. Let $\epsilon_{\theta}$ and $\epsilon_{\theta,\text{aug }}$ represent the noise predictions for the clear image I and augment image $\textbf{I}_{aug}$, respectively. Once these noises are identical and accurate, the final denoising depth will demonstrate consistency and robustness. Hence, as shown in Fig. 2(c), we reformulate $L_{ddim}$ as a novel noise-level trinity loss:
@@ 3.4. Multi-level Contrast
Figure 4. Visual results of each level trinity. Compared to the origin image, as shown in the red dashed rectangle, the image level trinity can assist in handling water surface artifacts and ground snow. The feature level trinity builds a coarse depth map with some wrong edges but the noise trinity fixes them all. Better viewed when zooming in.
@@
Feature level. We first extend it to the depth feature level, which exploits the robust ability of feature network $\mathcal{F}_{f}$. However, at this time, we lack the guidance as perfect as the added noise. Instead, as depicted in Fig. 2 (b), we adopt a teacher model $\mathcal{F}_{fT}$ to get the suboptimal guiding label and build the feature-level ‘trinity’ loss as:
(15) $\begin{array}[]{l}L_{feat}=\omega_{1}\|\mathcal{F}_{fS}(\textbf{I})-\mathcal{F}_{fS}(\textbf{I}_{aug})\|_{1}+\\ \omega_{2}(\|\mathcal{F}_{fT}(\textbf{I})-\mathcal{F}_{fS}(\textbf{I})\|_{1}+\|\mathcal{F}_{fT}(\textbf{I})-\mathcal{F}_{fS}(\textbf{I}_{aug})\|_{1}),\end{array}$
@@
(17) $L_{img}=\delta_{1}(\|\textbf{I}_{\text{aug }}^{\prime}-\textbf{I}\|_{1}+\|\textbf{I}^{\prime}-\textbf{I}\|_{1})+\delta_{2}\|\textbf{I}_{\text{aug }}^{\prime}-\textbf{I}^{\prime}\|_{1},$
where the values of $\delta_{1}$ and $\delta_{2}$ are same as those of $\omega_{1}$ and $\omega_{2}$.
@@
Table 1. Quantitative results on WeatherKITTI and KITTI dataset. For the error-based metrics, the lower value is better; and for the accuracy-based metrics, the higher value is better. The best and second-best results are marked in bold and underline.
@@
Figure 5. Qualitative results for the WeatherKITTI dataset. We compare D4RD with the current SoTA RMDE methods in the adverse rain and snow subsets. The part marked with ‘Clear’ is the corresponding sunny image (processed for more clarity). Regions with prominent differences are highlighted using dashed boxes.
@@ 3.5. Two stage training
The second training phase. The well-trained model from the first stage, now equipped with fixed weights, is designated as $\mathcal{F}_{T}$ for the second stage. We train $\mathcal{F}_{S}$ from scratch and further add $L_{feat}$ and $L_{img}$ to build the multi-level contrasts as mentioned in Section 3.4. The total loss of the second stage can be written as:
(19) $L_{stage2}=L_{nis}+L_{dis}+L_{ph}+\rho L_{e}+\eta(L_{feat}+L_{img}),$
where $\eta$ is empirically set to 0.5.
Figure 6. Qualitative results for the real weather dataset. Besides the RMDE methods, we also compare D4RD with the standard method MonoViT(Zhao et al., 2022b) in snowy and rainy environments to show our method’s practicality. The part marked with ’Magnification’ is processed for more clarity.
@@ 4.1. Dataset
WeatherKITTI(Wang et al., 2024) is built on the KITTI dataset and includes 6 kinds of weather augmentation. We train D4RD on this dataset because, compared with other synthetic datasets(Saunders et al., 2023; Mao et al., 2024), it is more realistic in simulating real-world complex scenarios. WeatherKITTI shares the same train/validate/test splits as KITTI. Therefore, for training, we follow Zhou’s split(Zhou et al., 2017) (containing 39,810 training images and 44,24 validation images), and for testing, we adopt Eigen’s split(Eigen et al., 2014) with 697 test images.
KITTI-C(Kong et al., 2023) is a synthetic dataset built for the RoboDepth Competition. In total, the benchmark contains 18 types of perturbations with 5 levels of corruption magnitudes, most of which are stronger than WeatherKITTI. We use it to fairly compare D4RD with their competition champions(Song et al., 2023) to further demonstrate our robust performance.
@@
Dense(Bijelic et al., 2020) is one of the few real snowy datasets and it has approximately 1,500 snowy images in total. We introduce this benchmark to test snowy scenes. To ensure consistency with the above datasets, we followed (Wang et al., 2024) and sampled 1 image every 3 to obtain 500 images as test data. To generate accurate depth GT in challenging snowy conditions, we use the DROR(Charron et al., 2018) algorithm to remove inaccurate depth readings from LiDARs affected by snowflakes. In addition, a small portion of distorted areas has been trimmed, resulting in a final resolution of $1880\times 924$. We will explain the reasons for not using the snowy part of the CADC dataset in the supplementary materials.
@@ 4.2. Implementation Details
D4RD is implemented in PyTorch and trained on a single NVIDIA RTX 3090. We use the AdamW optimizer(Loshchilov and Hutter, 2019) and other similar hyperparameters as MonoViT. The batch size is set to 6, and the training process runs 30 epochs in each stage. Considering the resolution differences across the datasets, we resize them to $640\times 192$ for both training and testing. For the diffusion process, we follow MonoDiffusion(Shao et al., 2023) and set it with 1000 training steps and 20 inference steps.
For fair comparisons, we retrained Robust-Depth(Saunders et al., 2023) and EC-Depth(Song et al., 2023) on the WeatherKITTI dataset. Furthermore, to demonstrate D4RD’s ability to handle robust scenes, we also retrain our model with the same training data used by (Song et al., 2023), named D4RD, to compare the performance on the KITTI-C dataset.
@@ 4.3. Quantitative and Qualitative Comparison
Here, we adopt the same evaluation metrics and errors up to 80m for all comparisons as in (Godard et al., 2019). Please refer to the supplementary materials for more details. The comparative methods include the baseline MonoDiffusion, traditional RMDE models’ baseline MonoViT(Zhao et al., 2022b), previous SoTA RMDE models WeatherDepth(Saunders et al., 2023), EC-Depth(Song et al., 2023), Robust-Depth(Saunders et al., 2023), and their retrained models. Besides, (Saunders et al., 2023) and (Wang et al., 2024) apply their approaches to various baselines, we want to clarify that we choose the one that uses MonoViT(Zhao et al., 2022b). In (Song et al., 2023), they take a two-stage training pipeline and we select the second stage model for the comparison.
@@
Table 4. Comparison of computation and time usage.
Model FLOPs(G) Param(M) Memory(GB) Train/Infer(s/batch) AbsRel
Monovit[36] 715.4 27.8 15.3 0.64/0.14 0.120
D4RD 881.9 27.9 23.5 1.02/0.32 0.099